OpenStudy (anonymous):

6-7(3-m) Help please

OpenStudy (anonymous):

6-7(3-m) = 6-21+7m = -15+7m = 7m -15

OpenStudy (anonymous):

Solving for m I assume?

OpenStudy (anonymous):

6-21-7m = 0? -15 -7m = 0 -15 = 7m -15/7 = m